OpenStudy (anonymous):

help me with this v v v The vertices of triangle IJK have coordinates I(−5, −2), J(−5, −4), and K(−4, −4). Which coordinates are the coordinates of the vertices of triangle IꞌJꞌKꞌ when triangle IJK is rotated 180° clockwise around the origin? A. Iꞌ(5, 2), Jꞌ(5, 4), Kꞌ(4, 4) B. Iꞌ(−2, 5), Jꞌ(−4, 5), Kꞌ(−4, 4) C. Iꞌ(2, −5), Jꞌ(4, −5), Kꞌ(4, −4) D. Iꞌ(−5, 2), Jꞌ(−5, 4), Kꞌ(−4, 4)

OpenStudy (anonymous):

@amistre64 @fallenangelorchid @mathstudent55 @bohotness

OpenStudy (bohotness):

okay what do you hink is your answer love

OpenStudy (anonymous):

i think its b

OpenStudy (bohotness):

okay why?

OpenStudy (anonymous):

if O is origin .....let P (x , y) with respect to axes OX and OY.....and let P(x' , y') with respect to axes OX' and OY' where angle X'OX = angle Y'OY = θ then x' = x cos(theta) + y sin(theta) y' = -x sin(theta) + y cos(theta)

OpenStudy (anonymous):

x' = -5 cos(180) + (-2) sin(180) y' = 5 sin(180) + (-2) cos(180)

OpenStudy (bohotness):

oay

OpenStudy (anonymous):

thats all i know XD

OpenStudy (bohotness):

see i if O is origin .....let P (x , y) with respect to axes OX and OY.....and let P(x' , y') with respect to axes OX' and OY' where angle X'OX = angle Y'OY = [Math Processing Error] then x' = x cos(theta) + y sin(theta) y' = -x sin(theta) + y cos(theta)

OpenStudy (bohotness):

u got the angle theta = 180 degrees.... so just plug the values

OpenStudy (bohotness):

like for I I' (x' , y') x' = -5 cos(180) + (-2) sin(180) y' = 5 sin(180) + (-2) cos(180)

OpenStudy (anonymous):

D?? :)

OpenStudy (bohotness):

does that information i give you sound like d?

OpenStudy (anonymous):

okay love, calm down. C?

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@cutiemarklove

OpenStudy (anonymous):

i would say it was D

OpenStudy (bohotness):

"_\\:)

OpenStudy (anonymous):

thanks dudeee

OpenStudy (anonymous):

^~^